# Tram: A Token-level Retrieval-augmented Mechanism for Source Code Summarization

### v2

@@ Abstract
Tong Ye1, Lingfei Wu2, Tengfei Ma3, Xuhong Zhang1, Yangkai Du1 Peiyu Liu1, Shouling Ji1, Wenhai Wang1thanks:  Corresponding author., 1Zhejiang University; 2Anytime.AI; 3Stony Brook University {tongye,zhangxuhong,yangkaidu,liupeiyu,sji,zdzzlab}@zju.edu.cn lwu@anytime-ai.com, tengfei.ma@stonybrook.edu
@@ 1 Introduction
These limitations have inspired us to explore a more fine-grained and sufficient retrieval approach on the summary generation process. In order to achieve the purpose of retrieving semantic similar summary tokens on the decoder side, we first construct a datastore to store the summary tokens and corresponding representations through a pre-trained base model offline. Meanwhile, to overcome the challenge of not fully utilizing code semantics on the encoder side when retrieving on the decoder side, we intelligently fuse summary token representation with code token representation and AST node representation with attention weight. This approach fully considers contextual code semantics associated with summary tokens. Then, at each generation step, the fused summary token representation is used to retrieve the top-$K$ most similar tokens. As illustrated in Table 1, the token-level retrieval results at the next token generation step “cos” are “cos, tangent, sin, hyperbolic, normal-$\cdots$. The retrieved top-$K$ tokens are expanded to a probability distribution, which we refer to as the retrieval-based distribution. The retrieval-based distribution is then fused with the vanilla distribution to form the final distribution. Additionally, our proposed token-level retrieval mechanism can be seamlessly integrated with existing sentence-level retrieval methods and code-related large pre-trained models.
@@ 3.6 Additional Sentence-level Retrieval
Datasets Java Python CCSD ${}^{\ddagger}$
@@ 4.1 Experimental Setup
We conduct the experiments on four public benchmarks of Java (Hu et al., 2018), Python (Wan et al., 2018), CCSD (C Code Summarization Dataset) (Liu et al., 2021), and ${}^{\ddagger}$ (Zhang et al., 2020). The partitioning of train/validation/test sets follows the original datasets. The statistics of the four datasets are shown in Table 2.
@@ 4.2 Baselines
Model CCSD ${}^{\ddagger}$
@@
Model Java ${}^{\ddagger}$
@@
Table 5: Human Evaluation on Java and ${}^{\ddagger}$ datasets.
@@ 4.3 Main Results
The main experiment results are shown in Table 3 and Table 4 in terms of three automatic evaluation metrics. The reason we have two tables is that transformer-based works compare their performance on the widely-used Java and Python benchmarks, while the retrieval-based works use two different benchmarks, namely CCSD and ${}^{\ddagger}$. Thus, our experiments are performed on all four datasets for a more thorough comparison. We calculate the metric values following the same scripts [^2] https://github.com/gingasan/sit3/blob/main/c2nl/eval/bleu/google_bleu.py.
@@
In Table 4, we compare Tram with other retrieval-based models on CCSD and ${}^{\ddagger}$ benchmarks. Our base model is even superior to other retrieval-based methods; the main reason is that the backbone  [^3] Other retrieval-based methods are RNN-based. are different. We reproduce Rencos architecture [^4] HGNN code is not open source. in our base model for a fair comparison, which we denoted as “Base + Rencos”. Tram outperforms all other retrieval-based methods, further improving performance with 2.05 BLEU points and 1.47 BLEU points on CCSD and ${}^{\ddagger}$, respectively. Furthermore, as shown in Table 3 and 4, enhancing Tram with additional sentence-level retrieval (refer as "Tram with SenRe") and its integration with code pre-trained models ("Our Method on Pre-trained Models" section in Table 3) leads to a notable improvement in performance.
@@ 4.4 Ablation Study
To validate the effectiveness of intelligently fusing summary token representation with code token representation $H_{t}$ and AST node representation $R_{t}$, we conduct an ablation experiment where we eliminate the $H_{t}$, $R_{t}$, and directly use $d_{t-1}$ to represent target summary token $s_{t}$ for comparison (refer as “Tram w/o HR”). As shown in Table 3 and 4, the performance declined by 0.47, 0.60, 0.21, and 0.32 BLEU points for Java, Python, CCSD, and ${}^{\ddagger}$, respectively. This decline in performance across all datasets demonstrated the importance of fusing code semantics into the summary token for effective token-level retrieval on the decoder side.
@@ 5.3 Datastore Quality and Robustness Analysis
To accurately assess the impact of datastore quality on Tram’s performance, we conduct robustness experiments where noise is intentionally introduced into the datastore. Specifically, we randomly shuffle a certain percentage of (representation, token) pairs, leading to misaligned pairings. These experiments, conducted using Python and Java datasets, are based on the averages from five separate runs. We introduce noise levels of 5%, 10%, and 20%, corresponding to the proportion of misaligned pairs in the datastore. The Table 8 presents the experimental results, indicating that even with a 10% noise level in the datastore, the BLEU score reduction is only up to 0.3 points. Furthermore, even under 20% noise conditions, the model maintains robust performance. These results suggest that the impact of datastore quality and the presence of noisy or poorly aligned pairs is relatively minimal, confirming the robustness of both the datastore and our Tram method.

### v3

@@ Abstract
Tong Ye1, Lingfei Wu2, Tengfei Ma3, Xuhong Zhang1, Yangkai Du1, Peiyu Liu1, Shouling Ji1, Wenhai Wang1thanks:  Corresponding author. 1Zhejiang University; 2Anytime.AI; 3Stony Brook University {tongye,zhangxuhong,yangkaidu,liupeiyu,sji,zdzzlab}@zju.edu.cn lwu@anytime-ai.com, tengfei.ma@stonybrook.edu
@@ 1 Introduction
These limitations have inspired us to explore a more fine-grained and sufficient retrieval approach on the summary generation process. In order to achieve the purpose of retrieving semantic similar summary tokens on the decoder side, we first construct a datastore to store the summary tokens and corresponding representations through a pre-trained base model offline. Meanwhile, to overcome the challenge of not fully utilizing code semantics on the encoder side when retrieving on the decoder side, we intelligently fuse summary token representation with code token representation and AST node representation with attention weight. This approach fully considers contextual code semantics associated with summary tokens. Then, at each generation step, the fused summary token representation is used to retrieve the top-$K$ most similar tokens. As illustrated in Table 1, the token-level retrieval results at the next token generation step “cos” are “cos, tangent, sin, hyperbolic, $\cdots$. The retrieved top-$K$ tokens are expanded to a probability distribution, which we refer to as the retrieval-based distribution. The retrieval-based distribution is then fused with the vanilla distribution to form the final distribution. Additionally, our proposed token-level retrieval mechanism can be seamlessly integrated with existing sentence-level retrieval methods and code-related large pre-trained models.
@@ 3.6 Additional Sentence-level Retrieval
Datasets Java Python CCSD Python
@@ 4.1 Experimental Setup
We conduct the experiments on four public benchmarks of Java (Hu et al., 2018), Python (Wan et al., 2018), CCSD (C Code Summarization Dataset) (Liu et al., 2021), and Python (Zhang et al., 2020). The partitioning of train/validation/test sets follows the original datasets. The statistics of the four datasets are shown in Table 2.
@@ 4.2 Baselines
Model CCSD Python
@@
Model Java Python
@@
Table 5: Human Evaluation on Java and Python datasets.
@@ 4.3 Main Results
The main experiment results are shown in Table 3 and Table 4 in terms of three automatic evaluation metrics. The reason we have two tables is that transformer-based works compare their performance on the widely-used Java and Python benchmarks, while the retrieval-based works use two different benchmarks, namely CCSD and Python. Thus, our experiments are performed on all four datasets for a more thorough comparison. We calculate the metric values following the same scripts [^2] https://github.com/gingasan/sit3/blob/main/c2nl/eval/bleu/google_bleu.py.
@@
In Table 4, we compare Tram with other retrieval-based models on CCSD and Python benchmarks. Our base model is even superior to other retrieval-based methods; the main reason is that the backbone  [^3] Other retrieval-based methods are RNN-based. are different. We reproduce Rencos architecture [^4] HGNN code is not open source. in our base model for a fair comparison, which we denoted as “Base + Rencos”. Tram outperforms all other retrieval-based methods, further improving performance with 2.05 BLEU points and 1.47 BLEU points on CCSD and Python, respectively. Furthermore, as shown in Table 3 and 4, enhancing Tram with additional sentence-level retrieval (refer as "Tram with SenRe") and its integration with code pre-trained models ("Our Method on Pre-trained Models" section in Table 3) leads to a notable improvement in performance.
@@ 4.4 Ablation Study
To validate the effectiveness of intelligently fusing summary token representation with code token representation $H_{t}$ and AST node representation $R_{t}$, we conduct an ablation experiment where we eliminate the $H_{t}$, $R_{t}$, and directly use $d_{t-1}$ to represent target summary token $s_{t}$ for comparison (refer as “Tram w/o HR”). As shown in Table 3 and 4, the performance declined by 0.47, 0.60, 0.21, and 0.32 BLEU points for Java, Python, CCSD, and Python, respectively. This decline in performance across all datasets demonstrated the importance of fusing code semantics into the summary token for effective token-level retrieval on the decoder side.
@@ 5.3 Datastore Quality and Robustness Analysis
To accurately assess the impact of datastore quality on Tram’s performance, we conduct robustness experiments where noise is intentionally introduced into the datastore. Specifically, we randomly shuffle a certain percentage of (representation, token) pairs, leading to misaligned pairings. These experiments, conducted using Python and Java datasets, are based on the averages from five separate runs. We introduce noise levels of 5%, 10%, and 20%, corresponding to the proportion of misaligned pairs in the datastore. Table 8 presents the experimental results, indicating that even with a 10% noise level in the datastore, the BLEU score reduction is only up to 0.3 points. Furthermore, even under 20% noise conditions, the model maintains robust performance. These results suggest that the impact of datastore quality and the presence of noisy or poorly aligned pairs is relatively minimal, confirming the robustness of both the datastore and our Tram method.
@@ 6 Conclusion
Acknowledgements
This work was partly supported by NSFC under Grant No. 62302443, the Fellowship of China National Postdoctoral Program for Innovative Talents (BX20230307), the Fundamental Research Funds for the Central Universities (Zhejiang University NGICS Platform). This research was also supported by the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.